# Statistical Test for Data Analysis Pipeline by Selective Inference

### v1

@@ 4.3 Update Rules
$\mathtt{update\_interval}(\mathcal{P},X,\bm{a},\bm{b},r)=([L_{z},U_{z}],\mathcal{M}_{\bm{a}+\bm{b}z},\mathcal{O}_{\bm{a}+\bm{b}z}).$ (22)
@@ 5 Implementations: Auto-Conditioning
As an example, Listing 1 shows a code example that defines the pipeline shown in Figure 2 based on our package pipelineprocessor and performs hypothesis testing on it. A similarly simple UI allows for easy implementation of other pipeline structures as well as automatic pipeline construction based on the cross-validation. For more examples, please refer to the Appendix G and the supplementary material.
Listing 1: Code example that defines the pipeline shown in Figure 2. We can implement an instance of the pipeline with inference methods simply by specifying each operation in turn. To perform hypothesis testing, we can call the inference method of the pipeline instance with the input dataset $(X,\bm{y})$ and the deviation of the noise $\sigma$.
@@
4def option1():
5 X, y = plp.make_dataset()
6 y = plp.mean_value_imputation(X, y)
@@
8 O = plp.soft_ipod(X, y, 0.02)
9 X, y = plp.remove_outliers(X, y, O)
@@
11 M = plp.marginal_screening(X, y, 5)
12 X = plp.extract_features(X, M)
@@
14 M1 = plp.stepwise_feature_selection(X, y, 3)
15 M2 = plp.lasso(X, y, 0.08)
16 M = plp.union(M1, M2)
17 return plp.make_pipeline(output=M)
@@
21M, p_list = pl.inference(X, y, sigma=1.0)
@@ 6 Numerical Experiments
In our experiments, we consider the three types of pipelines: op1, op2, and all_cv. The op1 and op2 are defined in Figure 2. The all_cv is a pipeline selected based on cross-validation from 16 different parameters sets each in the op1 and op2 pipelines (i.e., from 32 pipelines in total). For each three types of pipelines, we compare the proposed method (proposed) with oc (simple extension of SI literature to our setting) and naive test (naive), in terms of type I error rate and power. See Appendix D.1 for more details.
@@
In all experiments, we set the significance level $\alpha=0.05$. For the experiments to see the type I error rate, we change the number of samples $n\in\{100,200,300,400\}$ and set the number of features $d$ to $20$. See Appendix D.2 for the case when changing the number of features. In each setting, we generated 10,000 null datasets $(X,\bm{y})$, where $X_{ij}\sim\mathcal{N}(0,1),\forall(i,j)\in[n]\times[d]$, $\bm{y}\sim\mathcal{N}(0,I_{n})$ and each $y_{i},\forall i\in[n]$ was set to NaN with probability $0.03$. To investigate the power, we set $n=200$ and $d=20$ and generated 10,000 datasets $(X,\bm{y})$, where $X_{ij}\sim\mathcal{N}(0,1),\forall(i,j)\in[n]\times[d]$, $\bm{y}=X\bm{\beta}+\bm{\epsilon}$ with $\bm{\epsilon}\sim\mathcal{N}(0,I_{n})$, $\bm{\beta}\in\mathbb{R}^{d}$ is a vector whose first three elements are $\Delta$’s and the others are 00’s and each $y_{i},\ \forall i\in[n]$ was set to NaN with probability $0.03$. We set $\Delta\in\{0.2,0.4,0.6,0.8\}$. See Appendix D.3 for the computer resources used in the experiments.
@@ Results.
The results of type I error rate are shown in Figure 3. The proposed and oc successfully controlled the type I error rate under the significance level in all settings for all types of pipelines, whereas the naive could not. Because the naive failed to control the type I error rate, we no longer consider its power. The results of power are shown in Figure 4 and we confirmed that the proposed has much higher power than the oc in all settings for all types of pipelines.
@@
We compared the proposed method and the oc method for all_cv pipeline on eight real datasets: Energy Efficiency (heating load; Data1 and cooling load; Data2), Gas Turbine CO and NOx Emission Data Set (Data3), Wine Quality (red wine; Data4 and white wine; Data5), Abalone (Data6), Concrete Compressive Strength Data Set (Data7), and Real Estate Valuation (Data8) in the UCI Machine Learning Repository. From each of the original dataset, we randomly generated 1,000 sub-sampled datasets with size $n\in\{100,150,200\}$. We applied the proposed and oc to check the powers of the two methods. The results are shown in Table 1. In all datasets for all $n\in\{100,150,200\}$, the proposed method has much higher power than the oc method. Moreover, the power of our proposed method increases as $n$ increases.
(a) op1
(b) op2
(c) all_cv
Figure 3: Type I Error Rate when changing the number of samples $n$. Only our proposed method and the oc method are able to control the type I error rate in all settings for all types of pipelines.
(a) op1
(b) op2
(c) all_cv
Figure 4: Power when changing the true coefficient. Our proposed method has much higher power than the oc method in all settings for all types of pipelines.
@@
Data1 Data2 Data3 Data4 Data5 Data6 Data7 Data8
$n=100$59/.11 .54/.10 .74/.11 .33/.05 .33/.06 .34/.06 .54/.07 .50/.05
$n=150$70/.13 .60/.12 .79/.12 .43/.07 .38/.06 .43/.06 .69/.06 .63/.06
$n=200$75/.16 .69/.14 .82/.15 .46/.05 .45/.07 .55/.07 .74/.09 .77/.06
7 Conclusions, Limitations and Future Works
In this study, we introduced a novel framework for testing the statistical significance of data analysis pipelines based on the concept of SI. We believe that this study will contribute to improving the reliability of data-driven scientific research in various application fields. Currently, there are limitations on the applicable data analysis components, leaving several challenges in applying the proposed method to more complex data analysis pipelines. Furthermore, it is still challenging to select the optimal pipeline from all possible candidates, which requires exponentially increasing computational costs in our current implementation scheme.
@@ D.2 Additional Type I Error Rate Results
D.3 Computer Resources
@@ F.1 Statistical Inference after Cross-Validation
As a formulation of statistical inference after cross-validation, the discussion in §2 and §3 can be done in exactly the same way, with the only difference being the procedure for computing $\mathcal{M}$ and $\mathcal{O}$ (in §2 and §3, $\mathcal{M}$ and $\mathcal{O}$ are simply the outputs of a given mapping $\mathcal{P}$ representing a target pipeline). This implies that the procedure for computing the truncation intervals $\mathcal{Z}$ in §4 can not be directly applied to the case of cross-validation.
@@
We consider the case where $K$-fold cross-validation is performed. Let $(X,\bm{y})$ be the observed data set and $\{(T_{k},V_{k})\}_{k\in[K]}$ be the $K$ types of partition of training and validation sets, which satisfies $T_{k},V_{k}\in 2^{[n]}$, $T_{k}\cap V_{k}=\emptyset$, and $T_{k}\cup V_{k}=[n]$ for any $k\in[K]$. Then, the cross-validation error $E_{s}(X,y)$ for the pipeline $\mathcal{P}_{s}$ is defined as
$E_{s}(X,y)=\sum_{k\in[K]}\frac{1}{|V_{k}|}\|(D_{X}^{s}\bm{y})_{V_{k}}-X_{V_{k},\mathcal{M}_{s,k}}\hat{\bm{\beta}}_{s,k}(\bm{y})\|_{2}^{2},$ (61)
where $D_{X}^{s}$ is the linear transformation matrix in the missing value imputation of the pipeline $\mathcal{P}_{s}$, $\hat{\bm{\beta}}_{s,k}(\bm{y})=\left(X_{T_{k}\setminus\mathcal{O}_{s,k},\mathcal{M}_{s,k}}^{\top}X_{T_{k}\setminus\mathcal{O}_{s,k},\mathcal{M}_{s,k}}\right)^{-1}X_{T_{k}\setminus\mathcal{O}_{s,k},\mathcal{M}_{s,k}}^{\top}(D_{X}^{s}\bm{y})_{T_{k}\setminus\mathcal{O}_{s,k}}$, and $(\mathcal{M}_{s,k},\mathcal{O}_{s,k})$ is the output of the pipeline $\mathcal{P}_{s}$ with input $(X_{T_{k}},(D_{X}^{s}\bm{y})_{T_{k}})$. In $K$-fold cross-validation, the pipeline $\mathcal{P}_{s^{\ast}}$ is selected to minimize the cross-validation error $E_{s}(X,y)$, i.e., $s^{\ast}=\operatorname*{arg\,min}_{s\in[S]}E_{s}(X,y)$.
@@ F.2 Auto-Conditioning for Cross-Validation
$\operatorname*{arg\,min}_{s\in[S]}E_{s}(X,\bm{a}+\bm{b}r)=\operatorname*{arg\,min}_{s\in[S]}E_{s}(X,\bm{a}+\bm{b}z)(\coloneqq s^{\ast}),$ (63)
$\mathcal{P}_{s^{\ast}}(X,\bm{a}+\bm{b}r)=\mathcal{P}_{s^{\ast}}(X,\bm{a}+\bm{b}z).$ (64)
If we have this procedure, for any $r\in[L_{z},U_{z}]$, the selected features and the detected outliers after selecting the pipeline by cross-validation from the data set $(X,\bm{a}+\bm{b}r)$ are invariant. Therefore, the $p_{\mathrm{selective}}$ can be computed in exactly the same way as in §4.2. Hereafter, we provide the above procedure by extending Algorithm 1.
@@
$\forall r\in[L_{z}^{\mathrm{cv}},U_{z}^{\mathrm{cv}}],\ \operatorname*{arg\,min}_{s\in[S]}E_{s}(X,\bm{a}+\bm{b}r)=\operatorname*{arg\,min}_{s\in[S]}E_{s}(X,\bm{a}+\bm{b}z)(\coloneqq s^{\ast}),$ (65)
$\forall r\in[L_{z}^{\mathrm{sel}},U_{z}^{\mathrm{sel}}],\ \mathcal{P}_{s^{\ast}}(X,\bm{a}+\bm{b}r)=\mathcal{P}_{s^{\ast}}(X,\bm{a}+\bm{b}z),$ (66)
@@
$E_{s^{\ast}}(X,\bm{a}+\bm{b}r)\leq E_{s}(X,\bm{a}+\bm{b}r),\forall s\in[S],$ (68)
with $s^{\ast}=\operatorname*{arg\,min}_{s\in[S]}E_{s}(X,\bm{a}+\bm{b}z)$ become simultaneous quadratic inequalities, which can be solved analytically to finally obtain the interval $[L_{z}^{\mathrm{cv}},U_{z}^{\mathrm{cv}}]$.
To compute the interval $[L_{z}^{\mathrm{sel}},U_{z}^{\mathrm{sel}}]$, we simply use Algorithm 1 with input $(\mathcal{P}_{s^{\ast}},X,\bm{a},\bm{b},z)$.
@@ Appendix G Examples of Implementations
Listing 2: Code example that defines the pipeline referred to as op2 in the experiments. We can implement an instance of the pipeline with inference methods simply by specifying each operation in turn. To perform hypothesis testing, we can call the inference method of the pipeline instance with the input dataset $(X,\bm{y})$ and the deviation of the noise $\sigma$.
@@
4def option2():
5 X, y = plp.make_dataset()
6 y = plp.definite_regression_imputation(X, y)
@@
8 M = plp.marginal_screening(X, y, 5)
9 X = plp.extract_features(X, M)

### v2

@@ 4.3 Update Rules
$\mathtt{UpdateInterval}(\mathcal{P},X,\bm{a},\bm{b},r)\\ =([L_{z},U_{z}],\mathcal{M}_{\bm{a}+\bm{b}z},\mathcal{O}_{\bm{a}+\bm{b}z}).$ (27)
@@ 5 Implementations: Auto-Conditioning
As an example, Listing 1 shows a code example that defines the pipeline shown in Figure 2 based on our package si4automl and performs hypothesis testing on it. A similarly simple UI allows for easy implementation of other pipeline structures as well as automatic pipeline construction based on the cross-validation. For more examples, please refer to the Appendix G and the supplementary material.
Listing 1: Code example that defines the pipeline shown in Figure 2. We can create an instance of manager class which handles the desired pipeline simply by specifying each operation in turn. To perform hypothesis testing, we can call the inference method of the manager instance with the input dataset $(X,\bm{y})$ and the deviation of the noise $\sigma$.
@@
4def option1() -> PipelineManager:
5 X, y = initialize_dataset()
6 y = mean_value_imputation(X, y)
@@ 5 X, y = initialize_dataset()
8 O = soft_ipod(X, y, 0.02)
9 X, y = remove_outliers(X, y, O)
@@
11 M = marginal_screening(X, y, 5)
12 X = extract_features(X, M)
@@
14 M1 = stepwise_feature_selection(X, y, 3)
15 M2 = lasso(X, y, 0.08)
16 M = union(M1, M2)
17 return construct_pipelines(output=M)
@@ 16 M = union(M1, M2)
20X = np.random.normal(size=(100, 10))
21y = np.random.normal(size=100)
22M, p_list = pl.inference(X, y, sigma=1.0)
@@ 6 Numerical Experiments
In our experiments, we consider the three types of pipelines: op1, op2, and all_cv. The op1 and op2 are defined in Figure 2. The all_cv is a pipeline selected based on cross-validation from 16 different parameters sets each in the op1 and op2 pipelines (i.e., from 32 pipelines in total). For each three types of pipelines, we compare the proposed method (proposed) with oc (simple extension of SI literature to our setting) and naive test (naive), in terms of type I error rate and power. See Appendix D.1 for more details on the methods for comparison.
@@
In all experiments, we set the significance level $\alpha=0.05$. For the experiments to see the type I error rate, we change the number of samples $n\in\{100,200,300,400\}$ and set the number of features $d$ to $20$. See Appendix D.2 for the case when changing the number of features. In each setting, we generated 10,000 null datasets $(X,\bm{y})$, where $X_{ij}\sim\mathcal{N}(0,1),\forall(i,j)\in[n]\times[d]$, $\bm{y}\sim\mathcal{N}(0,I_{n})$ and each $y_{i},\forall i\in[n]$ was set to NaN with probability $0.03$. To investigate the power, we set $n=200$ and $d=20$ and generated 10,000 datasets $(X,\bm{y})$, where $X_{ij}\sim\mathcal{N}(0,1),\forall(i,j)\in[n]\times[d]$, $\bm{y}=X\bm{\beta}+\bm{\epsilon}$ with $\bm{\epsilon}\sim\mathcal{N}(0,I_{n})$, $\bm{\beta}\in\mathbb{R}^{d}$ is a vector whose first three elements are $\Delta$’s and the others are 00’s and each $y_{i},\ \forall i\in[n]$ was set to NaN with probability $0.03$. We set $\Delta\in\{0.2,0.4,0.6,0.8\}$. In addition, we provide an analysis of the computational time of the proposed method for larger datasets and more complex pipelines in Appendix D.3. See Appendix D.4 for the computer resources used in the experiments.
@@ Results.
The results of type I error rate are shown in left side of Figure 3. The proposed and oc successfully controlled the type I error rate under the significance level in all settings for all types of pipelines, whereas the naive could not. Because the naive failed to control the type I error rate, we no longer consider its power. The results of power are shown in right side of Figure 3 and we confirmed that the proposed has much higher power than the oc in all settings for all types of pipelines.
(a) Type I Error Rate and Power of op1 pipeline
(b) Type I Error Rate and Power of op2 pipeline
(c) Type I Error Rate and Power of all_cv pipeline
Figure 3: Type I Error Rate when changing the number of samples $n$ (left side) and Power when changing the true coefficient (right side). Both our proposed method and the oc method are able to control the type I error rate. However, our proposed method exhibits considerably higher power than the oc method in all settings for all types of pipelines.
@@
We compared the proposed and oc for all_cv pipeline on eight real datasets from the UCI Machine Learning Repository (all licensed under the CC BY 4.0; see Appendix D.5 for more details). From each of the original dataset, we randomly generated 1,000 sub-sampled datasets with size $n\in\{100,150,200\}$. We applied the proposed and oc to check the powers of the two methods. The results are shown in Table 1. In all datasets for all $n\in\{100,150,200\}$, the proposed has much higher power than the oc. Moreover, the power of our proposed method increases as $n$ increases.
@@
$n$ Data1 Data2 Data3 Data4 Data5 Data6 Data7 Data8
$100$57/.07 .48/.06 .57/.07 .51/.07 .68/.10 .55/.04 .30/.05 .25/.06
$150$79/.09 .71/.08 .66/.12 .57/.10 .74/.12 .72/.06 .37/.06 .37/.06
$200$91/.11 .80/.08 .78/.15 .66/.12 .76/.13 .82/.08 .49/.07 .40/.06
7 Conclusions
In this study, we introduced a novel framework for testing the statistical significance of feature selection pipelines comprising multiple MVI, OD, and FS algorithms based on the concept of SI. Our long-term goal extends beyond this current study to ensure the reproducibility of data-driven decision-making by accounting for the entire pipeline from raw data to final results, with this study on a class of feature selection pipelines serving as a proof of concept. To achieve this future goal, there are still limitations on the applicable data analysis components, presenting several challenges in extending the proposed framework to more complex data analysis pipelines. Additionally, it is interesting to consider extending this framework to scenarios where data analysis pipelines are automatically constructed using state-of-the-art AutoML approaches.
@@ D.2 Additional Type I Error Rate Results
D.3 Computational Time of the Proposed Method
We conducted experiments to investigate the computational time of our proposed method by applying it to three types of pipeline structures (Default, Parallel, and Serial) with large-scale datasets. Default pipeline correspond to the op1 pipeline in §6. Parallel and Serial pipelines are defined as in Figure 6, which clarifies the difference from the Default with components colored in pink. In the experiments, we change the number of samples $n\in\{400,800,1200,1600\}$ with the number of features $d=80$ and the number of features $d\in\{40,80,120,160\}$ with the number of samples $n=800$ to generate the null datasets in the same way as in the main experiments (§6). The results are shown in Figure 5.
(a) Number of Samples
(b) Number of Features
Figure 5: Computational Time when changing the number of samples and features. The results show that computational time is exponentially increased as the number of samples increases while the number of features has no obvious effect. Moreover, it can be seen that increasing the number of components in the pipeline increases the computational time, but how much it increases also depends on the structure.
Figure 6: Definition of Parallel and Serial pipelines used in Figure 5.
D.4 Computer Resources
@@
D.5 Details of the Real Datasets
We used the following eight real datasets from the UCI Machine Learning Repository. All datasets are licensed under the CC BY 4.0 license.
  • Airfoil Self-Noise (Brooks et al., 1989) for Data1
  • Concrete Compressive Strength (Yeh, 1998) for Data2
  • Energy Efficiency (Tsanas and Xifara, 2012) for Data3 (heating load) and Data4 (cooling load)
  • Gas Turbine CO and NOx Emission Data Set (gas, 2019) for Data5
  • Real Estate Valuation (Yeh, 2018) for Data6
  • Wine Quality (Cortez et al., 2009) for Data7 (red wine) and Data8 (white wine)
@@ F.1 Statistical Inference after Cross-Validation
As a formulation of statistical inference after cross-validation, the discussion in §2 and §3 can be done in exactly the same way, except with two changes: (i) the procedure for computing $\mathcal{M}$ and $\mathcal{O}$ (in §2 and §3, $\mathcal{M}$ and $\mathcal{O}$ are simply the outputs of a given mapping $\mathcal{P}$ representing a target pipeline), and (ii) the dependence on the response vector $\bm{y}$ of which method to use for missing value imputation. This implies that the procedure for computing the truncation intervals $\mathcal{Z}$ in §4 can not be directly applied to the case of cross-validation.
@@
We consider the case where $K$-fold cross-validation is performed. Let $(X,\bm{y})$ be the observed data set and $\{(T_{k},V_{k})\}_{k\in[K]}$ be the $K$ types of partition of training and validation sets, which satisfies $T_{k},V_{k}\in 2^{[n]}$, $T_{k}\cap V_{k}=\emptyset$, and $T_{k}\cup V_{k}=[n]$ for any $k\in[K]$. Then, the cross-validation error $E_{s}(X,\bm{y})$ for the pipeline $\mathcal{P}_{s}$ is defined as
$E_{s}(X,\bm{y})=\sum_{k\in[K]}\frac{1}{|V_{k}|}\|(D_{X}^{s}\bm{y})_{V_{k}}-X_{V_{k},\mathcal{M}_{s,k}}\hat{\bm{\beta}}_{s,k}(\bm{y})\|_{2}^{2},$ (66)
where $D_{X}^{s}$ is the linear transformation matrix in the missing value imputation of the pipeline $\mathcal{P}_{s}$, $\hat{\bm{\beta}}_{s,k}(\bm{y})=\left(X_{T_{k}\setminus\mathcal{O}_{s,k},\mathcal{M}_{s,k}}^{\top}X_{T_{k}\setminus\mathcal{O}_{s,k},\mathcal{M}_{s,k}}\right)^{-1}X_{T_{k}\setminus\mathcal{O}_{s,k},\mathcal{M}_{s,k}}^{\top}(D_{X}^{s}\bm{y})_{T_{k}\setminus\mathcal{O}_{s,k}}$, and $(\mathcal{M}_{s,k},\mathcal{O}_{s,k})$ is the output of the pipeline $\mathcal{P}_{s}$ with input $(X_{T_{k}},(D_{X}^{s}\bm{y})_{T_{k}})$. In $K$-fold cross-validation, the pipeline $\mathcal{P}_{s^{\ast}}$ is selected to minimize the cross-validation error $E_{s}(X,\bm{y})$, i.e., $s^{\ast}=\operatorname*{arg\,min}_{s\in[S]}E_{s}(X,y)$.
@@ F.2 Auto-Conditioning for Cross-Validation
$\operatorname*{arg\,min}_{s\in[S]}E_{s}(X,\bm{a}+\bm{b}r)=\operatorname*{arg\,min}_{s\in[S]}E_{s}(X,\bm{a}+\bm{b}z)(\coloneqq s(z)),$ (68)
$\mathcal{P}_{s(z)}(X,\bm{a}+\bm{b}r)=\mathcal{P}_{s(z)}(X,\bm{a}+\bm{b}z).$ (69)
If we have this procedure, for any $r\in[L_{z},U_{z}]$, the selected features and the detected outliers after selecting the pipeline by cross-validation from the data set $(X,\bm{a}+\bm{b}r)$ are invariant. Therefore, the $p_{\mathrm{selective}}$ can be computed in exactly the same way as in §4 only by adding the condition $D_{X}^{s(z)}=D_{X}^{s^{\ast}}$ as well as the condition $\mathcal{M}_{\bm{a}+\bm{b}z}=\mathcal{M}_{\bm{y}}$ and $\mathcal{O}_{\bm{a}+\bm{b}z}=\mathcal{O}_{\bm{y}}$. Hereafter, we provide the above procedure by extending Algorithm 1.
@@
$\forall r\in[L_{z}^{\mathrm{cv}},U_{z}^{\mathrm{cv}}],\ \operatorname*{arg\,min}_{s\in[S]}E_{s}(X,\bm{a}+\bm{b}r)=\operatorname*{arg\,min}_{s\in[S]}E_{s}(X,\bm{a}+\bm{b}z)(\coloneqq s(z)),$ (70)
$\forall r\in[L_{z}^{\mathrm{sel}},U_{z}^{\mathrm{sel}}],\ \mathcal{P}_{s(z)}(X,\bm{a}+\bm{b}r)=\mathcal{P}_{s(z)}(X,\bm{a}+\bm{b}z),$ (71)
@@
$E_{s(z)}(X,\bm{a}+\bm{b}r)\leq E_{s}(X,\bm{a}+\bm{b}r),\forall s\in[S],$ (73)
with $s(z)=\operatorname*{arg\,min}_{s\in[S]}E_{s}(X,\bm{a}+\bm{b}z)$ become simultaneous quadratic inequalities, which can be solved analytically to finally obtain the interval $[L_{z}^{\mathrm{cv}},U_{z}^{\mathrm{cv}}]$.
To compute the interval $[L_{z}^{\mathrm{sel}},U_{z}^{\mathrm{sel}}]$, we simply use Algorithm 1 with input $(\mathcal{P}_{s(z)},X,\bm{a},\bm{b},z)$.
@@ Appendix G Examples of Implementations
Listing 2: Code example that defines the pipeline referred to as op2 in the experiments. We create an instance of manager class which handles the pipeline simply by specifying each operation in turn. To perform hypothesis testing, we can call the inference method of the manager instance with the input dataset $(X,\bm{y})$ and the deviation of the noise $\sigma$.
@@
4def option2() -> PipelineManager:
5 X, y = initialize_dataset()
6 y = definite_regression_imputation(X, y)
@@ 5 X, y = initialize_dataset()
8 M = marginal_screening(X, y, 5)
9 X = extract_features(X, M)